# An Objective Improvement Approach

### v2

@@ 6. Experimental Evaluation
To evaluate the performance of the novel approach, we have implemented it in an experimental framework in C++. Together with the objective improvement algorithm (OI, for short) we have also implemented the classic asymmetric variant of the strategy improvement algorithm (SI, for short) that has been described in Section 3. Note that our implementation uses floating point numbers with double precision (type doubles), while the approach itself assumes infinite precision. As a consequence, match checks are not possible and the solutions can be affected by an $\epsilon$ error due to the precision of the library (that comes from the external linear programming solver too). For most parts, this could be rectified by using fractions. However, while fractions with infinite precision are available with the right libraries, their use would lead to an overhead and would not solve the precision problem of the external linear programming solver. We also note that the small noise added in Section 5.4 to almost surely make a game sharp and improving requires a continuous value space.
@@
One cost of doing this is that we do not have a similarly straightforward integration of (3) non-local updates, where we are in the optimal solution for an objective function that cannot be improved with rule (2). Instead of using the method from Theorem 5 to slightly adjust a simplex method to find an improvement for improving games (knowing that the games we create are almost surely improving), we use the general rule from Lemma 4. In principle, this might lead to large plateaux that are hard to exit, but we considered this very unlikely, at least for the improving games we create, and our experiments have confirmed this.
@@
The number of iterations (Figure 2) represents how many times the linear programming solver has been called. As expected, SI is more efficient than OI at solving these games in terms of LP calls. What was less expected is how the advantage of SI over OI grows with the size of the game. Solving the linear programs, instead, is simpler for OI, at least when using a simplex style algorithm [^6] Simplex style algorithms (reported in Algorithm 1) operate in two phases: in the first phase, they look for any basis that defines a solution to the constraint system (Line 1), and in the second, they optimise (while loop). For OI, the first phase can be skipped for all but the first call; this is because the constraint system does not change, so that the previous solution can be used as a starting point. For SI, the constraint system changes; here, the previous solution is never a solution to the new linear program.. We also note that only 55 (5.5%) of the 1000 games considered overall reached a non-improving state. In these rare cases, this has been solved by choosing a stale switch (i.e. a switch of strategy that leads to the same valuation value).
@@ 7. Discussion
Our results for random games with a low (five to ten) outdegree suggests that the number of local updates seen grows linearly with the size of the game for both SI and OI, and that the number of iterations initially grows and later plateaus. This may well be the same for games with a minimal outdegree of two, and the data clearly supports this for SI, but the number of LP calls growth almost linearly for OI. This might be an outlier in the behaviour and it is well possible that OI plateaus later, as it is unlikely that games with a low and very low outdegree behave fundamentally different. For SI, this was entirely expected, and it is maybe not overly surprising for OI either. Generally speaking, OI appears to do better than SI measured in the number of LP calls on random games, except where the outdegree is tiny.
@@
Seeing that (1) and (2) are equally cheap and simple, we could also favour (2) over (1): where we have a solution $\mathsf{val}$ defined by a basis, we can first find a joint strategy $\sigma$ such that $f_{\sigma}(\mathsf{val})$ is minimal among all strategies, before turning to (1) and (3). This would also make sure that no vertex is ever visited twice. This could happen in principle when favouring (1) over (2), as a step that is good for one objective function might not be good for its successor, but as the value of the objective function goes down in every step, a return to a previously optimal valuation is not possible.

### v3

@@ 6. Experimental Evaluation
To evaluate the performance of the novel approach, we have implemented it in an experimental framework in C++. Together with the objective improvement algorithm (OI, for short), we have also implemented the classic asymmetric variant of the strategy improvement algorithm (SI, for short) that has been described in Section 3. Note that our implementation uses floating-point numbers with double precision (type double), while the approach itself assumes infinite precision. As a consequence, match checks are not possible and the solutions can be affected by an $\epsilon$ error due to the precision of the library (that comes from the external linear programming solver too). For most parts, this could be rectified by using fractions. However, while fractions with infinite precision are available with the right libraries, their use would lead to an overhead and would not solve the precision problem of the external linear programming solver. We also note that the small noise added in Section 5.4 to almost surely make a game sharp and improving requires a continuous value space.
One cost of doing this is that we do not have a similarly straightforward integration of (3) non-local updates, where we are in the optimal solution for an objective function that cannot be improved with rule (2). Instead of using the method from Theorem 5 to slightly adjust a simplex method to find an improvement for improving games (knowing that the games we create are almost surely improving), we use the general rule from Lemma 4, trying random stale strategies. In principle, this might lead to large plateaux that are hard to exit, but we considered this very unlikely, at least for the improving games we create, and our experiments have confirmed this.
@@
The number of iterations (Figure 2) represents how many times the linear programming solver has been called. As expected, SI is more efficient than OI at solving these games in terms of LP calls. What was less expected is how the advantage of SI over OI grows with the size of the game. Solving the linear programs, instead, is simpler for OI, at least when using a simplex style algorithm [^9] Simplex style algorithms (reported in Algorithm 1) operate in two phases: in the first phase, they look for any basis that defines a solution to the constraint system (Line 1), and in the second, they optimise (while loop). For OI, the first phase can be skipped for all but the first call; this is because the constraint system does not change, so that the previous solution can be used as a starting point. For SI, the constraint system changes; here, the previous solution is never a solution to the new linear program.. We also note that only 55 (5.5%) of the 1000 games considered overall reached a non-improving game. In these rare cases, this has been solved by choosing a stale switch (i.e. a switch of strategy that leads to the same valuation value).
@@ 7. Discussion
Our results for random games with a low (five to ten) outdegree suggests that the number of local updates seen grows linearly with the size of the game for both SI and OI, and that the number of iterations initially grows and later plateaus. This may well be the same for games with a minimal outdegree of two, and the data clearly supports this for SI, but the number of LP calls growth almost linearly for OI. This might be an outlier in the behaviour and it is well possible that OI plateaus later, as it is unlikely that games with a low and very low outdegree behave fundamentally different. For SI, this was entirely expected, and it is perhaps not overly surprising for OI either. Generally speaking, OI appears to do better than SI measured in the number of LP calls on random games, except where the outdegree is tiny.
@@
Seeing that (1) and (2) are equally cheap and simple, we could also favour (2) over (1): where we have a solution $\mathsf{val}$ defined by a basis, we can first find a joint strategy $\sigma$ such that $f_{\sigma}(\mathsf{val})$ is minimal among all strategies, before turning to (1) and (3). This would also make sure that no corner of the polytope is ever visited twice. This could happen in principle when favouring (1) over (2), as a step that is good for one objective function might not be good for its successor, but as the value of the objective function goes down in every step, a return to a previously optimal valuation is not possible.